MATH 132, section 1 – Calculus for Physical and Life Sciences Discussion 3 – Table of Integrals Practice February 12, 2008

Do all the integrals below by finding an appropriate entry from the table in the text, and put the work for your group’s integral up on the board. Note: Some of these will require preliminary algebra and/or a u-substitution to match a table entry exactly.

1. (I) and (V)

∫ x^{2} sin(x^{3}) cos(3x^{3}) dx

2. (II) and (VI)

∫ \frac{dx}{(3 + 5e^{x})^{2}}

3. (III) and (VII)

∫ \frac{dx}{x\sqrt{4(ln(x))^{2} + 4 ln(x) + 2}}

4. (IV)

∫ \frac{sin(x)dx}{\sqrt{4 cos^{2}(x) + 4 cos(x) − 3}}